Whispering Winds had total debt of $507,500 and $760,350 as at December 31, 2020, and December 31, 2019, respectively, of which $105,000 and $150,000 was current. In addition, the company had total assets of $875,000 and $685,000 as at December 31, 2020, and December 31, 2019, respectively, of which $130,000 and $165,000 was current. Evaluate the company’s debt-paying ability in 2020 and 2019.

2020 2019

Debt-paying ability by calculating the debt to total assets ratio: (Round answers to 0 decimal places, e.g. 500%.)

Debt-paying ability by calculating current ratio: (Round answers to 2 decimal places, e.g. 45.50.)
Answer #1

Debt to Assets Ratio

=Total liabilities/Total Assets.

December 2019

=( $ 760,350/685000 ) x 100 = 111%

Which means 111% of companies assets are financed by debt.

December 2020

= ( 507,500 /875,000 ) x 100 = 58%

Which means 58% of total Assets of companies is financed by debt.

2.Current Ratio

= Current Assets/ current liabilities

December 2019

= (165,000/150,000 )= 1.10

December 2020

= (130,000/105,000) = 1.24